# Epidemiology of Large Language Models: A Benchmark for Observational Distribution Knowledge

### v1

@@ 1 Introduction
Artificial intelligence (AI) systems hold great potential to accelerate many scientific disciplines including healthcare (jiang2017artificial; shaheen2021applications), education (holmes2022state), and economics (bickley2022artificial). Automated systems based on AI are also increasingly used in a wide variety of real-world settings. Yet, despite rapid progress and broad adoption of these systems, many core capabilities required for robust and reliable AI still need to be developed further, on the path to reaching artificial human intelligence (bubeck2023sparks; kaplan2020scaling). One of the hallmark features expected from AI systems is the ability to reason probabilistically about the world, and not only to mimic human intuition and language. Such capabilities are needed for systems capable of assisting scientific discovery, advancing our understanding of complex phenomena, and help societal decision-making.
@@
In this work, our focus is on observational probabilistic knowledge, where the term observational refers to the first layer of Pearl’s Causal Hierarchy (PCH, for short) pearl:2k; bareinboim2020on. The PCH distinguishes different types of probabilistic knowledge, corresponding to different cognitive capabilities of (AI) systems: Layer 1 of observation (seeing), which includes reasoning about observed events and correlations between variables; Layer 2 of intervention (doing), which includes inferences about systems in which causal interventions are performed; and Layer 3 of counterfactual reasoning (imagining, see Fig. 2 left), which allows inference for hypothesized events for which the required pre-conditions may not have materialized in reality. Important results from previous work demonstrate that, in absence of appropriate assumptions, moving across different layers of PCH is provably impossible. The result known as the Causal Hierarchy Theorem (CHT, (bareinboim2020on)) states that observational knowledge (Layer 1) about a system underdetermines its interventional (Layer 2) and counterfactual (Layer 3) behavior. With causal modeling and assumptions, inferences across layers may become feasible. A classical example is causal modeling that uses observational data, combined with structural assumptions, to infer interventional or counterfactual distributions pearl:2k; bareinboim2016causal. Therefore, if one can show that LLMs do not have access to observational knowledge (Layer 1), based on the CHT, one may also be skeptical of these models’ capabilities for reliable statements about interventions or counterfactuals.
@@
While LLMs have been tremendously successful at various tasks, some caution is nonetheless warranted. Classical results in statistics highlight fundamental limitations of learning distributions in higher dimensions: the famous curse of dimensionality results demonstrate theoretically that learning distributions becomes exponentially harder (in terms of required samples) as dimensionality grows. Seminal work by Charles Stone stone1982optimal establishes that estimation rates degrade sharply with dimensionality, with the optimal rate of convergence for an unknown differentiable regression function $f$ being $\mathcal{O}(n^{-\frac{1}{2+d}})$ in any $L^{p}$ norm, where $d$ is the dimension of the input of $f$. Such results challenge the notion of universal distributional learning in higher dimensions, in stark contrast with the notion of LLMs working as universal approximators.
@@ 1.1 Contributions
A preview of our results is shown in Fig. 1, indicating that over a range of models, the performance of the current generation of LLMs in terms of knowledge of real-world distributions is quite poor. As we argue later in the text, access to observational distributions is a fundamental step towards performing interventional (Layer 2) or counterfactual (Layer 3) inference, implying that are our results also implicitly challenge the abilities of LLMs in higher layers of Pearl’s Hierarchy.
@@ 2 Benchmark Axes
We use the language of structural causal models (SCMs) (pearl:2k). An SCM is a tuple $\mathcal{M}:=\langle V,U,\mathcal{F},P(u)\rangle$ , where $V$, $U$ are sets of endogenous (observable) and exogenous (latent) variables, respectively, $\mathcal{F}$ is a set of functions $f_{V_{i}}$, one for each $V_{i}\in V$, where $V_{i}\leftarrow f_{V_{i}}(\mathrm{pa}(V_{i}),U_{V_{i}})$ for some $\mathrm{pa}(V_{i})\subseteq V$ and $U_{V_{i}}\subseteq U$. The assignment mechanisms $\mathcal{F}$ determine how each variable attains its value, and the set $\mathrm{pa}(V_{i})$ is called the parent set of $V_{i}$. Together with the probability distribution $P(u)$ over the exogenous variables $U$, the SCM specifies the entire behavior of the underlying phenomenon, meaning that it fully specifies all observational, interventional, and counterfactual distributions (bareinboim2020on).
The remainder of the section is organized according to Fig. 2, and the different axes appearing therein. We first discuss observational knowledge and its different dimensionalities (Sec. 2.1), followed by the description of datasets that are used (Sec. 2.2), and then the tasks constructed for eliciting models’ capabilities (Sec. 2.2.1). In Sec. 2.3, we discuss different models currently included in the benchmark.
2.1 Axis 1: Dimensionality of Layer 1 (Observational) Knowledge
Our focus in this paper is on evaluating observational (Layer 1) knowlege available in AI systems. The observational distribution of the underlying phenomenon can be defined through the SCM itself:
Definition 1 (Observational Distribution (bareinboim2020on)).
An SCM $\mathcal{M}$ that is a 4-tuple $\langle V,U,\mathcal{F},P(u)\rangle$ induces a joint probability distribution $P(V)$ such that for each $Y\subseteq V$,
$P^{\mathcal{M}}(Y=y)=\sum_{u}\mathbb{1}\Big(Y(u)=y\Big)P(U=u),$ (1)
where $Y(u)$ is the solution for $Y$ after evaluating $\mathcal{F}$ with $U=u$.
In the above definition, $Y\subseteq V$ should be thought of as multidimensional. A fixed value of exogenous $U=u$ is called a unit, and can be thought of as a individual in the population. To obtain the observational distribution $P(Y=y)$, we go over all units $u$, and add up the probability mass $P(U=u)$ of all units for which $Y$ attains the value $y$. The observational distribution $P(V)$ captures all probabilistic relationships among observed variables under passive observation. On top of the distribution, we can also consider other types of quantities, such as (conditional) moments. In particular, we may consider different types of knowledge, such as:
@@ 2.1 Axis 1: Dimensionality of Layer 1 (Observational) Knowledge
The above types of knowledge can be viewed as a hierarchy, with the level of required knowledge being progressively more refined. Our goal can be described as follows: we design tasks probing the above types of knowledge, based on large-scale datasets describing populations. For instance, we refer to a dataset $\mathcal{D}$, and consider the distribution over covariates $P(V)$ implied by $\mathcal{D}$. Then, we elicit responses from an LLM about the same distribution. The model’s distribution is labeled $\tilde{P}(V)$. Our goal is to then compare various aspects of $\tilde{P}(V)$ and $P(V)$, according to types of knowledge (i)-(v). In this way, we can systematically evaluate whether LLMs internalize such quantities over real-world populations. In fact, we can formally argue that testing knowledge on observational distributions is a critical first step for understanding knowledge on causal capabilities more broadly:
Proposition 1 (Causal Hierarchy Theorem (bareinboim2020on)).
  1. (a)
  2. (b)
Corollary 2 (No Observational Distribution $\implies$ No Layer 2/3 Inference).
If a model’s distribution $\tilde{P}(V)$ differs from the true $P(V)$, no guarantees can be provided for the validity of the model’s interventional or counterfactual inferences.
The above corollary captures an important motivation for the benchmark constructed in this paper – even if we assume that an LLM has access to correct causal knowledge (a point we do not investigate in this manuscript), it still may not be able to perform causal inferences in case its observational distribution does not match the one in the real world. Therefore, testing whether language models have access to observational knowledge (required for causal reasoning) is an important first step since a negative answer warrants skepticism about the model’s capabilities for performing claims in higher layers of the PCH.
@@ 2.2 Axis 2: Datasets Describing Populations
As mentioned in the previous section, we will leverage large-scale datasets that describe population level statistics in order to establish the ground truth observational distribution $P(V)$. After this, we will probe the language model’s distribution $\tilde{P}(V)$ to draw comparisons. In this work, for the observational distributions $P(V)$, we make use of ten large, publicly available datasets that collectively describe diverse aspects of the US population, and are often considered representative of the national level statistics:
@@
    Department of Labor BLS Data 2023 (bls2023): The Bureau of Labor Statistics of the US Department of Labor provides detailed information about occupations, including worker demographics. We investigate occupations by sex and race.
@@
Each of the datasets either reports national-level statistics, or provides data on individuals. When individual data is available, sample weights are provided (determining how many persons in the overall US population the individual represents), allowing one to obtain representative results at the national level. Together, these datasets allow for a broad evaluation of models’ abilities to approximate population-level observational distributions across social, economic, educational, health, and behavioral domains. We next describe how different tasks in our benchmark are constructed.
@@ 2.2.1 Axis 2 Tasks: Responses and Scoring
Our benchmark consists of different tasks, and each task is associated with a dataset. The task is defined by the pair $V_{Y},V_{X}$, representing the conditional distribution $P(V_{Y}\mid V_{X})$ (here, $V_{X}$ could be possibly empty). Each task is also accompanied by a natural language prompt template $\pi$ (which takes a value of $V_{X}=v_{X}$ and returns a natural language question), and a set of valid answers in the domain of $V_{Y}$, labeled $\mathrm{dom}(V_{Y})$.
Example 1 (NSDUH: Marijuana Usage by Age).
@@
Intuitively, the difficulty of the task should increase with larger dimensions of $V_{Y}$ and $V_{X}$. For instance, if $|V_{Y}|=1$, $|V_{X}|=0$, our task is to recover the 1-dimensional marginal distribution $P(V_{Y})$. If $|V_{Y}|=1,|V_{X}|>0$, we are trying to recover a 1-dimensional conditional distribution $P(V_{Y}\mid V_{X})$. The question templates $\pi$ are carefully designed, to reflect the original questions asked to individuals during data acquisition. In total, we constructed 75 tasks across 10 datasets for $|V_{Y}|=1,|V_{X}|=1$ (we refer to this as the low-dimensional setting), and 94 tasks across 4 datasets for $|V_{Y}|=1,|V_{X}|\in\{2,3,4,5\}$ and $V_{Y}$ binary (referred to as the high-dimensional setting). Fig. 2 (bottom left) offers some further examples, in addition to Ex. 1, while Appendix A contains the full list of tasks. We remark here that an alternative way of eliciting the model’s knowledge also exists, captured by the following example:
Example 1 (continued – NSDUH: Marijuana Usage by Age).
For the task of eliciting the distribution of marijuana usage by age group, a template prompt
[…]
may be used instead, representing an alternative way of eliciting $\tilde{P}(V_{Y}\mid V_{X})$.
The comparison of the two prompting approaches (probabilistic vs. Q&A) is discussed in Appendix LABEL:appendix:qa-vs-prob, while in the main text we focus on Q&A prompt templates as in Eq. 2.
Eliciting Responses.
As mentioned, for each question the model is asked, there is a specific range of possible responses. Our approach to obtain the model answers is to label each response in $\mathrm{dom}(V_{Y})$ with labels $A,B,C,\dots$, and then inspect the conditional probability associated with tokens $A,B,C,$ etc. Let $\mathcal{A}(v_{Y})$ denote the letter of the answer $v_{Y}$. The model’s probability associated with each answer $v_{Y}\in\mathrm{dom}(V_{Y})$ corresponding to label $\mathcal{A}({v}_{Y})$ in the prompt is then simply computed as
@@
Example 1 (continued – NSDUH: Marijuana Usage by Age).
@@
This is a known strategy for eliciting model responses, used in numerous works on question answering (dominguez2024questioning; hendrycks2020measuring; santurkar2023whose). To mitigate ordering bias (dominguez2024questioning), we average over different answer permutations. For open-source models we use next-token probabilities; for closed models we use Monte Carlo bin-counting. The ground truth observational distribution $P(V_{Y}\mid V_{X})$ is estimated based on the available dataset, using bin-counting in low-dimensional settings and lightgbm plus cross-validation in high-dimensional settings. See Appendix LABEL:app:elicitation for more details on eliciting responses.
@@
We next develop a scoring strategy for each task based on the true distribution $P(V_{Y}\mid V_{X})$ and the model’s distribution $\tilde{P}(V_{Y}\mid V_{X})$. For this purpose, we use the $L_{1}$-norm (instead of KL divergence, to avoid sensitivity to low-probability events and support mismatch (gibbs2002choosing)), and define the following distributional distance:
@@
In this context, we refer to the $P^{\text{unif}}$ distribution as a random baseline. In the high-dimensional case, which focuses on binary $V_{Y}$, we add another baseline, namely a fixed 0/1 prediction depending on the dataset mean, given by $P^{0/1}(V_{Y}=1\mid V_{X}=v_{X})\overset{\Delta}{=}\mathbb{1}(\mathbbm{E}_{P}[V_{Y}]>0.5)$ for all $v_{Y},v_{X}$, where $\mathbbm{E}_{P}[V_{Y}]$ represent the true distribution mean. Therefore, the 0/1 baseline predicts a constant probability of 0 for all conditioning sets for outcomes with a marginal incidence $\mathbbm{E}_{P}[V_{Y}]\leq 0.5$, and predicts a constant $1$ whenever the marginal incidence is $\mathbbm{E}_{P}[V_{Y}]>0.5$. Such a baseline also corresponds to effectively no probabilistic knowledge. With these baselines in place, our scoring system is described next (illustrated in Fig. 3(a)). As many datasets consist of samples of the population (and do not survey the entire population), there is some uncertainty on the ground truth distribution $P$. In the scoring, we account for this fact, and proceed as follows. We draw bootstrap samples of size $|\mathcal{D}|$ (size of the dataset) labeled $\mathcal{D}^{(b)}$, and extract the bootstrap ground truth distribution $P^{(b)}$. We then look at the distance $D(P||P^{(b)})$ of $P$ (based on the available data) and $P^{(b)}$ (obtained from bootstrapped data), across different samples (see green density in Fig. 3(a)). The upper 5% quantile of the spread of $D(P||P^{(b)})$ is labeled with the perfect score $S=100$ – meaning that the score 100 is assigned if a model’s distribution cannot be statistically distinguished from the ground truth at the 5% significance level. For setting the score of 0, we use the minimum of the distances $D(P||P^{\text{unif}})$, $D(P||P^{0/1})$, with the latter only considered for binary outcomes. The model’s final score is then linearly interpolated between $S=0$ and S=100S={100}, as follows:
@@
[…]
(a) Scoring strategy.
(b) Model leaderboard.
Figure 3: Scoring strategy and model performance across task groups.

### v2

@@ Abstract
footnotetext: Project page available at https://llm-observatory.org/.
Keywords: generative artificial intelligence, large language models, probabilistic knowledge
@@ 1 Introduction
Artificial intelligence (AI) systems hold great potential to accelerate many scientific disciplines including healthcare (jiang2017artificial; shaheen2021applications), education (holmes2022state), and economics (bickley2022artificial), just to name a few. Automated systems based on AI are also increasingly used in a wide variety of real-world settings. Yet, despite rapid progress and broad adoption of these systems, many core capabilities required for robust and reliable AI still need to be developed further, on the path to reaching artificial human intelligence (bubeck2023sparks; kaplan2020scaling; hendrycks2025definition). One of the hallmark features expected from AI systems is the ability to reason probabilistically about the world, and not only to mimic human intuition and language. Such capabilities are needed for systems capable of assisting scientific discovery, advancing our understanding of complex phenomena, and help societal decision-making.
@@
In this work, our focus is on observational probabilistic knowledge, where the term observational refers to the first layer of Pearl’s Causal Hierarchy (PCH, for short) (pearl:2k; bareinboim2020on). The PCH distinguishes different types of probabilistic knowledge, corresponding to different cognitive capabilities of (AI) systems: Layer 1 of observation (seeing), which includes reasoning about observed events and correlations between variables; Layer 2 of intervention (doing), which includes inferences about systems in which causal interventions are performed; and Layer 3 of counterfactual reasoning (imagining, see Fig. 2 left), which allows inference for hypothesized events for which the required pre-conditions may not have materialized in reality. Important results from previous work demonstrate that, in absence of appropriate assumptions, moving across different layers of PCH is provably impossible. The result known as the Causal Hierarchy Theorem (CHT, (bareinboim2020on)) states that observational knowledge (Layer 1) about a system underdetermines its interventional (Layer 2) and counterfactual (Layer 3) behavior. With causal modeling and assumptions, inferences across layers may become feasible. A classical example is causal modeling that uses observational data, combined with structural assumptions, to infer interventional or counterfactual distributions (pearl:2k; bareinboim2016causal). Therefore, if one can show that LLMs do not have access to observational knowledge (Layer 1), based on the CHT, one may also be skeptical of these models’ capabilities for reliable statements about interventions or counterfactuals.
@@
While LLMs have been tremendously successful at various tasks, some caution is nonetheless warranted. Classical results in statistics highlight fundamental limitations of learning distributions in higher dimensions: the famous curse of dimensionality results demonstrate theoretically that learning distributions becomes exponentially harder (in terms of required samples) as dimensionality grows. Seminal work by stone1982optimal establishes that estimation rates degrade sharply with dimensionality, with the optimal rate of convergence for an unknown differentiable regression function $f$ being $\mathcal{O}(n^{-\frac{1}{2+d}})$ in any $L^{p}$ norm, where $d$ is the dimension of the input of $f$. Such results challenge the notion of universal distributional learning in higher dimensions, in stark contrast with the notion of LLMs working as universal approximators.
@@ 1.1 Contributions
A preview of our results is shown in Fig. 1, indicating that over a range of models, the performance of the current generation of LLMs in terms of knowledge of real-world distributions is quite poor. Through an example, we emphasize why this type of knowledge may be important for the end user:
Example 1 (Diabetes Rates)
A user with persistent fatigue is unsure whether their symptoms may be indicative of diabetes. They provide an LLM with basic demographic and lifestyle information, such as age, sex, BMI, smoking status, and ask: “Given my age (45), female sex, BMI (29), non-smoker status, what is the baseline probability that I have diabetes?”
Answering the query above reliably requires access to the observational distribution governing disease risk in the population, and our later investigation demonstrates that current models do not possess this knowledge. The illustrative example is just one instance of a query where observational knowledge is essential, and similar type of knowledge is relevant across many other domains investigated in this manuscript.
@@
Relationship to Causal Inference.
While not directly focused on causal knowledge, our investigation still bears important implications for causal inference, which we now elucidate. Following the standard approach in the literature (pearl:2k), a structural causal model (SCM) is a tuple $\mathcal{M}:=\langle V,U,\mathcal{F},P(u)\rangle$, where $V$, $U$ are sets of endogenous (observable) and exogenous (latent) variables, respectively, $\mathcal{F}$ is a set of functions $f_{V_{i}}$, one for each $V_{i}\in V$, where $V_{i}\leftarrow f_{V_{i}}(\mathrm{pa}(V_{i}),U_{V_{i}})$ for some $\mathrm{pa}(V_{i})\subseteq V$ and $U_{V_{i}}\subseteq U$. The assignment mechanisms $\mathcal{F}$ determine how each variable attains its value, and the set $\mathrm{pa}(V_{i})$ is called the parent set of $V_{i}$. Together with the probability distribution $P(u)$ over the exogenous variables $U$, the SCM specifies the entire behavior of the underlying phenomenon, meaning that it fully specifies all observational, interventional, and counterfactual distributions (bareinboim2020on).
The result known as Causal Hierarchy Theorem (bareinboim2020on) shows formally that access to observational distributions is a fundamental step towards performing interventional (Layer 2) or counterfactual (Layer 3) inference:
Proposition 1 (Causal Hierarchy Theorem (bareinboim2020on))
  1. (a)
  2. (b)
Corollary 2 (No Observational Distribution $\implies$ No Layer 2/3 Inference)
If the model’s distribution $\tilde{P}(V)$ differs from the true $P(V)$, no guarantees can be provided for the validity of the model’s interventional or counterfactual inferences.
The above corollary captures an important motivation for the benchmark constructed in this paper – even if we assume that an LLM has access to correct causal knowledge (a point we do not investigate in this manuscript), it still may not be able to perform causal inferences in case its observational distribution does not match the one in the real world. Interestingly, existing works in the literature treat language models as already possessing causal knowledge (de2025efficient), while our work provides evidence against such practice.
@@ 2 Benchmark Axes
This section is organized according to Fig. 2, and the different axes appearing therein. We first discuss (observational) probabilistic knowledge and its different dimensionalities (Sec. 2.1), followed by the description of datasets that are used (Sec. 2.2), and then the tasks constructed for eliciting models’ capabilities (Sec. 2.2.1). In Sec. 2.3, we discuss different models currently included in the benchmark.
2.1 Axis 1: Dimensionality of Probabilistic Knowledge
The observational distribution $P(V)$ over variables $V$ captures all probabilistic relationships among observed variables under passive observation. To test whether a model has access to the correct distribution, we may focus on different quantities, such as:
@@ 2.1 Axis 1: Dimensionality of Probabilistic Knowledge
The above types of knowledge can be viewed as a hierarchy, with the level of required knowledge being progressively more refined. Our goal can be described as follows: we design tasks probing the above types of knowledge, based on large-scale datasets describing populations. For instance, we refer to a dataset $\mathcal{D}$, and consider the distribution over covariates $P(V)$ implied by $\mathcal{D}$. Then, we elicit responses from an LLM about the same distribution. The model’s distribution is labeled $\tilde{P}(V)$. Our goal is to then compare various aspects of $\tilde{P}(V)$ and $P(V)$, according to types of knowledge (i)-(v). In this way, we can systematically evaluate whether LLMs internalize such quantities over real-world populations.
@@ 2.2 Axis 2: Datasets Describing Populations
We leverage large-scale datasets that describe population level statistics in order to establish the ground truth observational distribution $P(V)$. After this, we probe the language model’s distribution $\tilde{P}(V)$ to draw comparisons. In this work, for the observational distributions $P(V)$, we make use of ten large, publicly available datasets that collectively describe diverse aspects of the US population, and are considered representative of national level statistics:
@@
    Department of Labor BLS Data 2023 (bls2023): The Bureau of Labor Statistics of the US Department of Labor provides information about demographics across occupations. We investigate occupations by sex and race.
@@
Each of the datasets either reports national-level statistics, or provides data on individuals. When individual data is available, sample weights are provided (determining how many persons in the overall US population the individual represents), allowing one to obtain representative results at the national level. Together, these datasets allow us to (approximately) establish population-level observational distributions across social, economic, educational, health, and behavioral domains. We next describe how different tasks in our benchmark are constructed.
@@ 2.2.1 Axis 2 Tasks: Responses and Scoring
Our benchmark consists of different tasks, and each task is associated with a dataset. The task is defined by the pair $V_{Y},V_{X}$, representing the conditional distribution $P(V_{Y}\mid V_{X})$ (here, $V_{X}$ could be possibly empty). Each task is also accompanied by a natural language prompt template $\pi$ (which takes a value of $V_{X}=v_{X}$ and returns a natural language question), and a set of valid answers in the domain of $V_{Y}$, labeled $\mathrm{dom}(V_{Y})$. For each task, we consider two ways of eliciting model responses, referred to as (a) question-answer (QA) prompting and (b) likelihood prompting.
Question & Answer Prompting.
The first approach for eliciting the model’s distribution is QA prompting, illustrated in the following example:
Example 2 (NSDUH: Marijuana Usage by Age)
@@
Intuitively, the difficulty of the task should increase with larger dimensions of $V_{Y}$ and $V_{X}$. For instance, if $|V_{Y}|=1$, $|V_{X}|=0$, our task is to recover the 1-dimensional marginal distribution $P(V_{Y})$. If $|V_{Y}|=1,|V_{X}|>0$, we are trying to recover a 1-dimensional conditional distribution $P(V_{Y}\mid V_{X})$. The question templates $\pi$ are carefully designed, to reflect the original questions asked to individuals during data acquisition. In total, we constructed 75 tasks across 10 datasets for $|V_{Y}|=1,|V_{X}|=1$ (we refer to this as the low-dimensional setting), and 94 tasks across 4 datasets for $|V_{Y}|=1,|V_{X}|\in\{2,3,4,5\}$ and $V_{Y}$ binary (referred to as the high-dimensional setting). Fig. 2 (bottom left) offers some further examples, in addition to Ex. 2, while Appendix A contains the full list of tasks.
For QA prompting, our approach to obtain the model answers is to label each response in $\mathrm{dom}(V_{Y})$ with labels $A,B,C,\dots$, and then inspect the conditional probability associated with tokens $A,B,C,$ etc. Let $\mathcal{A}(v_{Y})$ denote the letter of the answer $v_{Y}$. The model’s probability associated with each answer $v_{Y}\in\mathrm{dom}(V_{Y})$ corresponding to label $\mathcal{A}({v}_{Y})$ in the prompt is then simply computed as
@@
Example 2 (continued – NSDUH: Marijuana Usage by Age)
@@
This is a known strategy for eliciting model responses, used in numerous works on question answering (dominguez2024questioning; hendrycks2020measuring; santurkar2023whose), and many of the large benchmarks in the literature (hendrycks2020measuring; clark2018think; bisk2020piqa). As models may exhibit ordering bias (dominguez2024questioning), we consider all of the different $|\mathrm{dom}(V_{Y})|!$ permutations of labels $A,B,C,\dots$, over the answers $v_{Y}\in\mathrm{dom}(V_{Y})$, and average the probabilities accordingly (if the number of permutations $|\mathrm{dom}(V_{Y})|!>120$, we consider 120 random permutations). The above way of eliciting responses may be used for all models with open-source weights, which allows us to access next-token prediction probabilities. For the ground truth distribution $P$, in the low-dimensional setting we use a simple bin-counting estimator on the available dataset, whereas in the high-dimensional case we fit the distribution $P(V_{Y}\mid V_{X})$ using lightgbm (ke2017lightgbm), with actual values $P(V_{Y}=v_{Y}\mid V_{X}=v_{x})$ obtained out-of-sample, using 5-fold cross-validation.
Likelihood Prompting.
An alternative way of eliciting the distribution $P(V_{Y}=v_{Y}\mid V_{X}=v_{x})$ from the model is to inquire about the probability directly:
Example 2 (continued – NSDUH: Marijuana Usage by Age)
For the task of eliciting the distribution of marijuana usage by age group, we use the following prompt:
What is the probability that a person aged 16 years ever used marijuana?
    A. 0%   B. 0%-5%   ...   U. 95%-100%   V. 100%
We then inspect the next-token prediction probabilities in $\tilde{P}^{\mathrm{nt}}$, and choose the response with the highest probability:
The chosen response letter $\mathcal{A}^{\text{resp}}$ is then mapped to either an interval, or a fixed value of 0% or 100%. For assigning the final predicted probability, for intervals, we take the midpoint. For instance, if the answer B. 0%-5% is chosen, we set the predicted probability to $0.025$.
For likelihood prompting, permutations over answer labels are not considered.
@@
After explaining how model answers are obtained, we next develop a scoring strategy for each task based on the true distribution $P(V_{Y}\mid V_{X})$ and the elicited model’s distribution $\tilde{P}(V_{Y}\mid V_{X})$. For this purpose, we use the $L_{1}$-norm (instead of KL divergence, to avoid sensitivity to low-probability events and support mismatch (gibbs2002choosing)), and define the following distributional distance:
@@
[…]
Figure 3: Scoring strategy.
In this context, we refer to the $P^{\text{unif}}$ distribution as a random baseline. In the high-dimensional case, which focuses on binary $V_{Y}$, we add another baseline, namely a fixed 0/1 prediction depending on the dataset mean, given by $P^{0/1}(V_{Y}=1\mid V_{X}=v_{X})\overset{\Delta}{=}\mathbb{1}(\mathbbm{E}_{P}[V_{Y}]>0.5)$ for all $v_{Y},v_{X}$, where $\mathbbm{E}_{P}[V_{Y}]$ represent the true distribution mean. Therefore, the 0/1 baseline predicts a constant probability of 0 for all conditioning sets for outcomes with a marginal incidence $\mathbbm{E}_{P}[V_{Y}]\leq 0.5$, and predicts a constant $1$ whenever the marginal incidence is $\mathbbm{E}_{P}[V_{Y}]>0.5$. Such a baseline also corresponds to effectively no probabilistic knowledge. With these baselines in place, our scoring system is described next (illustrated in Fig. 3). As many datasets consist of samples of the population (and do not survey the entire population), there is some uncertainty on the ground truth distribution $P$. In the scoring, we account for this fact, and proceed as follows. We draw bootstrap samples of size $|\mathcal{D}|$ (size of the dataset) labeled $\mathcal{D}^{(b)}$, and extract the bootstrap ground truth distribution $P^{(b)}$. We then look at the distance $D(P||P^{(b)})$ of $P$ (based on the available data) and $P^{(b)}$ (obtained from bootstrapped data), across different samples (see green density in Fig. 3). The upper 5% quantile of the spread of $D(P||P^{(b)})$ is labeled with the perfect score $S=100$ – meaning that the score 100 is assigned if a model’s distribution cannot be statistically distinguished from the ground truth at the 5% significance level. For setting the score of 0, we use the minimum of the distances $D(P||P^{\text{unif}})$, $D(P||P^{0/1})$, with the latter only considered for binary outcomes. The model’s final score is then linearly interpolated between $S=0$ and S=100S={100}, as follows: